OpenStudy (anonymous):

Does the experimental probability of an event equal its theoretical probability?

OpenStudy (anonymous):

only when the number of trials tends to be very high, theoretically number of trials should reach infinity then only theoretical and practical probablities becomes equal.

OpenStudy (anonymous):

so i have to Provide mathematical examples to support your opinions. You may use tables or diagrams to organize and present your thoughts. any ideas??

OpenStudy (anonymous):

U want me to support my opinion with data?

OpenStudy (anonymous):

if you can, that would be great i just dont really understand the whole thing

OpenStudy (anonymous):

Okay wait il provide u

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

theoretical probability- what should happen after testing experimental probability- what did happen after testing This is true, but to expand on this: Suppose we toss a fair (non weighted) coin. The theoretical probability of getting a head (or tail) is 1/2 because the coin is fair so we should get an equal number of heads and tails ie. this is what should happen in theory. In practice this doesn't always happen, which is where experimental probability comes in. Experimental probability is the probability of an outcome of an event based on an experiment. For example, if we toss the coin 10 times and get 4 heads and 6 tails we would say that the experimental probability of getting a head is 4/10 = 2/5 and the experimental probability of getting a tail is 6/10 = 3/5. The more experiments we do, the closer the probabilities get to the theoretical probability. Experimental probability is particularly useful when looking at problems which involved events and outcomes for which we don't know a theoretical probability, so we use experimental probability to get an approximation.